LISTSERV mailing list manager LISTSERV 16.0

Help for HEANET-FAULTS Archives


HEANET-FAULTS Archives

HEANET-FAULTS Archives


June 2008


View:

Show Author | Hide Author
Show Table of Contents | Hide Table of Contents

 

Table of Contents:

HEA-NOC/20070904-1 [UPDATE] modsecurity install on samhain.heanet.ie (3 messages)
HEA-NOC/20071210-9 [UPDATE] DIAS connections to new IP equipment (2 messages)
HEA-NOC/20080327-8 [CLOSE] Fault on Videoconferencing ISDN Gateway (1 message)
HEA-NOC/20080403-3 [CLOSE] 10G link LAN extension Install (1 message)
HEA-NOC/20080407-3 [CLOSE] Degraded service to TippInst, RCSI and NCIRL (1 message)
HEA-NOC/20080410-1 [CLOSE] Loss of connectivity to GTI (1 message)
HEA-NOC/20080410-5 [CLOSE] NUIG Network Transition Second Phase (1 message)
HEA-NOC/20080410-5 [UPDATE] NUIG Network Transition Second Phase (2 messages)
HEA-NOC/20080424-8 [CLOSE] Removing UCC Old Links (1 message)
HEA-NOC/20080513-1 [CLOSE] DCC fibre Cut (1 message)
HEA-NOC/20080519-2 [CLOSE] NUI Migration to New IP Network (1 message)
HEA-NOC/20080519-2 [UPDATE] NUI Migration to New IP Network (1 message)
HEA-NOC/20080521-20 [CLOSE] Primary 10GE Link to Geant via London Down (1 message)
HEA-NOC/20080524-1 [CLOSE] GEANT Primary 10GE to London is down (1 message)
HEA-NOC/20080527-10 [CLOSE] Problems viewing GTI streams from RTE Streaming LAN (1 message)
HEA-NOC/20080527-12 [CLOSE] Bringing up second BGP session to CAO on cr2-kp (1 message)
HEA-NOC/20080527-12 [UPDATE] Bringing up second BGP session to CAO on cr2-kp (2 messages)
HEA-NOC/20080527-6 [CLOSE] WIT PoP Switch Outages (1 message)
HEA-NOC/20080527-9 [CLOSE] e|net Planned Work (1 message)
HEA-NOC/20080528-2 [UPDATE] Carlow College Migration to New IP Network (1 message)
HEA-NOC/20080528-3 [UPDATE] Cessation of legacy Eircom ITnet links (3 messages)
HEA-NOC/20080529-4 [CLOSE] Citywest Core Router Upgrade (1 message)
HEA-NOC/20080529-4 [UPDATE] Citywest Core Router Upgrade (1 message)
HEA-NOC/20080601-1 [CLOSE] Geant link to London (1 message)
HEA-NOC/20080601-1 [UPDATE] Geant link to London (3 messages)
HEA-NOC/20080603-2 [OPEN] Planned maintenance in the NIRAN Pop in NWRC, Derry. (1 message)
HEA-NOC/20080603-3 [CLOSE] Setup bonded interface on samhain (1 message)
HEA-NOC/20080603-3 [OPEN] Setup bonded interface on samhain (1 message)
HEA-NOC/20080603-3 [UPDATE] Setup bonded interface on samhain (2 messages)
HEA-NOC/20080603-6 [CLOSE] Link Flaps on GEANT GENEVA Circuit (1 message)
HEA-NOC/20080603-6 [OPEN] Link Flaps on GEANT GENEVA Circuit (1 message)
HEA-NOC/20080604-3 [CLOSE] Dublin and National Ring Fibre Moves (1 message)
HEA-NOC/20080604-3 [OPEN] Dublin and National Ring Fibre Moves (1 message)
HEA-NOC/20080604-3 [UPDATE] Dublin and National Ring Fibre Moves (2 messages)
HEA-NOC/20080604-7 [CLOSE] Carlow College Outage (1 message)
HEA-NOC/20080604-7 [OPEN] Carlow College Outage (1 message)
HEA-NOC/20080605-6 [CLOSE] Mary Immaculate UL loss connectivity (1 message)
HEA-NOC/20080605-6 [OPEN] Mary Immaculate UL loss connectivity (1 message)
HEA-NOC/20080605-6 [UPDATE] Mary Immaculate UL loss connectivity (1 message)
HEA-NOC/20080606-2 [CLOSE] Maintenace Work - Outage notice 18/06/2008 (1 message)
HEA-NOC/20080606-2 [OPEN] Maintenace Work - Outage notice 18/06/2008 (1 message)
HEA-NOC/20080606-2 [UPDATE] Maintenace Work - Outage notice 18/06/2008 (1 message)
HEA-NOC/20080606-4 [CLOSE] Loss of connectivity Tallaght Hostpital (1 message)
HEA-NOC/20080606-4 [OPEN] Loss of connectivity Tallaght Hostpital (1 message)
HEA-NOC/20080606-6 [OPEN] IT Tralee connection move (1 message)
HEA-NOC/20080606-6 [UPDATE] IT Tralee connection move (1 message)
HEA-NOC/20080606-7 [OPEN] JANET connection recabling in Citywest (1 message)
HEA-NOC/20080606-7 [UPDATE] JANET connection recabling in Citywest (1 message)
HEA-NOC/20080606-8 [CLOSE] DKIT Connectivity issues (1 message)
HEA-NOC/20080606-8 [OPEN] DKIT Connectivity issues (1 message)
HEA-NOC/20080607-1 [CLOSE] Brief outage to St. Angela's College (1 message)
HEA-NOC/20080607-1 [OPEN] Brief outage to St. Angela's College (1 message)
HEA-NOC/20080607-1 [UPDATE] Brief outage to St. Angela's College (4 messages)
HEA-NOC/20080607-3 [CLOSE] Scheduled outage to Dark Fibre on NTFON (1 message)
HEA-NOC/20080607-3 [OPEN] Scheduled outage to Dark Fibre on NTFON (1 message)
HEA-NOC/20080607-4 [CLOSE] Alarms detected on BT circuit HEAN107523 (1 message)
HEA-NOC/20080607-4 [OPEN] Alarms detected on BT circuit HEAN107523 (1 message)
HEA-NOC/20080609-5 [OPEN] IADT Down (1 message)
HEA-NOC/20080609-5 [UPDATE] IADT Down (3 messages)
HEA-NOC/20080610-2 [OPEN] Routing policy change to General Internet (1 message)
HEA-NOC/20080610-2 [UPDATE] Routing policy change to General Internet (1 message)
HEA-NOC/20080610-7 [CLOSE] restart of Primary DNS server ns.heanet.ie (1 message)
HEA-NOC/20080610-7 [OPEN] restart of Primary DNS server ns.heanet.ie (1 message)
HEA-NOC/20080611-11 [CLOSE] Replacement of memory card in Drumline ADVA node. (1 message)
HEA-NOC/20080611-11 [OPEN] Replacement of memory card in Drumline ADVA node. (1 message)
HEA-NOC/20080611-15 [OPEN] An Chéim Migration (1 message)
HEA-NOC/20080611-15 [UPDATE] An Chéim Migration (2 messages)
HEA-NOC/20080611-16 [CLOSE] DC electrical work in CWT PoP (1 message)
HEA-NOC/20080611-16 [OPEN] DC electrical work in CWT PoP (1 message)
HEA-NOC/20080611-16 [UPDATE] DC electrical work in CWT PoP (1 message)
HEA-NOC/20080612-15 [OPEN] e-Net Planned Outage Notice (1 message)
HEA-NOC/20080612-15 [UPDATE] e-Net Planned Outage Notice (1 message)
HEA-NOC/20080612-4 [CLOSE] Removal of Old Redundant Connections to Old IP Network (1 message)
HEA-NOC/20080612-4 [OPEN] Removal of Old Redundant Connections to Old IP Network (1 message)
HEA-NOC/20080613-1 [OPEN] Tipp Inst Site outage (1 message)
HEA-NOC/20080613-1 [UPDATE] Tipp Inst Site outage (6 messages)
HEA-NOC/20080613-11 [CLOSE] No connectivity to GTI (1 message)
HEA-NOC/20080613-11 [OPEN] No connectivity to GTI (1 message)
HEA-NOC/20080613-11 [UPDATE] No connectivity to GTI (3 messages)
HEA-NOC/20080613-17 [CLOSE] Repatching of Global Crossing fibre for GEANT in Esat-X (1 message)
HEA-NOC/20080613-17 [OPEN] Repatching of Global Crossing fibre for GEANT in Esat-X (1 message)
HEA-NOC/20080613-17 [UPDATE] Repatching of Global Crossing fibre for GEANT in Esat-X (3 messages)
HEA-NOC/20080613-9 [OPEN/CLOSE] Link to GTI flapped (1 message)
HEA-NOC/20080616-2 [OPEN] ESB fibre work affecting Galway PoP (1 message)
HEA-NOC/20080616-4 [CLOSE] HEAnet NOC Phone System At Risk (1 message)
HEA-NOC/20080616-4 [OPEN] HEAnet NOC Phone System At Risk (1 message)
HEA-NOC/20080617-6 [OPEN] Multicast routing not working at IT Athlone (1 message)
HEA-NOC/20080617-8 [OPEN] TCD transition to new IP network (1 message)
HEA-NOC/20080617-8 [UPDATE] TCD transition to new IP network (3 messages)
HEA-NOC/20080618-6 [CLOSE] Migration of ns.heanet.ie to new hardware (1 message)
HEA-NOC/20080618-6 [OPEN] Migration of ns.heanet.ie to new hardware (1 message)
HEA-NOC/20080618-6 [UPDATE] Migration of ns.heanet.ie to new hardware (2 messages)
HEA-NOC/20080619-5 [CLOSE] IPv6 connectivity to WIT disrupted (1 message)
HEA-NOC/20080619-5 [OPEN] IPv6 connectivity to WIT disrupted (1 message)
HEA-NOC/20080619-5 [UPDATE] IPv6 connectivity to WIT disrupted (1 message)
HEA-NOC/20080621-1 [OPEN] The circuit to the Hermes Project schools is down (1 message)
HEA-NOC/20080621-1 [UPDATE] The circuit to the Hermes Project schools is down (1 message)
HEA-NOC/20080623-11 [CLOSE] UCD IP Link Transition (1 message)
HEA-NOC/20080623-11 [OPEN] UCD IP Link Transition (1 message)
HEA-NOC/20080623-11 [UPDATE] UCD IP Link Transition (1 message)
HEA-NOC/20080623-13 [OPEN] Removal of Old Circuits (1 message)
HEA-NOC/20080623-13 [UPDATE] Removal of Old Circuits (1 message)
HEA-NOC/20080623-9 [CLOSE] Repair work on fibres in the HEAnet Servecentric PoP (1 message)
HEA-NOC/20080623-9 [OPEN] Repair work on fibres in the HEAnet Servecentric PoP (1 message)
HEA-NOC/20080623-9 [UPDATE] Repair work on fibres in the HEAnet Servecentric PoP (3 messages)
HEA-NOC/20080624-1 [OPEN] An Cheim Failover Testing (1 message)
HEA-NOC/20080624-1 [UPDATE] An Cheim Failover Testing (1 message)
HEA-NOC/20080624-12 [OPEN] Luas Works affecting NCI backup circuit (1 message)
HEA-NOC/20080624-13 [OPEN] ISDN gateway service line maintenance (1 message)
HEA-NOC/20080624-3 [CLOSE] Network maintenance at IT Tralee (1 message)
HEA-NOC/20080624-3 [OPEN] Network maintenance at IT Tralee (1 message)
HEA-NOC/20080624-5 [CLOSE] Netflow Server Reboot Required (1 message)
HEA-NOC/20080624-5 [OPEN] Netflow Server Reboot Required (1 message)
HEA-NOC/20080624-7 [CLOSE] Serial backup circuit to NCRI down (1 message)
HEA-NOC/20080624-7 [OPEN] Serial backup circuit to NCRI down (1 message)
HEA-NOC/20080624-9 [OPEN] Scheduled Maintenance on JANET network 8th July 2008 (1 message)
HEA-NOC/20080624-9 [UPDATE] Scheduled Maintenance on JANET network 8th July 2008 (1 message)
HEA-NOC/20080626-8 [CLOSE] UL wish to cease using the hotstandby service. (1 message)
HEA-NOC/20080626-8 [OPEN] UL wish to cease using the hotstandby service. (1 message)
HEA-NOC/20080626-8 [UPDATE] UL wish to cease using the hotstandby service. (1 message)
HEA-NOC/20080627-1 [OPEN] IRCSET migration within HEA IT infrastructure. (1 message)
HEA-NOC/20080627-1 [UPDATE] IRCSET migration within HEA IT infrastructure. (2 messages)
HEA-NOC/20080630-1 [OPEN] BT Fibre Maintenance - Cork <-> Mallow 20080710 - 23:50 (1 message)

LISTSERV Archives

LISTSERV Archives

HEANET-FAULTS Home

HEANET-FAULTS Home

Subject Sorted by Subject, Chronologically

From

Date

Size

New Thread

HEA-NOC/20070904-1 [UPDATE] modsecurity install on samhain.heanet.ie

HEA-NOC/20070904-1 [UPDATE] modsecurity install on samhain.heanet.ie

Rob Gallagher

Wed, 11 Jun 2008 12:04:11 +0100

70 lines

HEA-NOC/20070904-1 [UPDATE] modsecurity install on samhain.heanet.ie

Rob Gallagher

Tue, 10 Jun 2008 17:42:00 +0100

63 lines

HEA-NOC/20070904-1 [UPDATE] modsecurity install on samhain.heanet.ie

Rob Gallagher

Tue, 10 Jun 2008 15:58:21 +0100

60 lines

New Thread

HEA-NOC/20071210-9 [UPDATE] DIAS connections to new IP equipment

HEA-NOC/20071210-9 [UPDATE] DIAS connections to new IP equipment

Dave Wilson

Thu, 12 Jun 2008 10:12:00 +0100

51 lines

HEA-NOC/20071210-9 [UPDATE] DIAS connections to new IP equipment

Dave Wilson

Wed, 11 Jun 2008 15:47:28 +0100

47 lines

New Thread

HEA-NOC/20080327-8 [CLOSE] Fault on Videoconferencing ISDN Gateway

HEA-NOC/20080327-8 [CLOSE] Fault on Videoconferencing ISDN Gateway

Kevin Dermody

Thu, 5 Jun 2008 09:39:40 +0100

47 lines

New Thread

HEA-NOC/20080403-3 [CLOSE] 10G link LAN extension Install

HEA-NOC/20080403-3 [CLOSE] 10G link LAN extension Install

Cormac O'Cianain

Tue, 24 Jun 2008 09:14:56 +0100

44 lines

New Thread

HEA-NOC/20080407-3 [CLOSE] Degraded service to TippInst, RCSI and NCIRL

HEA-NOC/20080407-3 [CLOSE] Degraded service to TippInst, RCSI and NCIRL

Garreth Malone

Tue, 24 Jun 2008 15:34:52 +0100

47 lines

New Thread

HEA-NOC/20080410-1 [CLOSE] Loss of connectivity to GTI

HEA-NOC/20080410-1 [CLOSE] Loss of connectivity to GTI

Garreth Malone

Tue, 24 Jun 2008 15:33:00 +0100

45 lines

New Thread

HEA-NOC/20080410-5 [CLOSE] NUIG Network Transition Second Phase

HEA-NOC/20080410-5 [CLOSE] NUIG Network Transition Second Phase

Brian Nisbet

Fri, 27 Jun 2008 10:04:39 +0100

57 lines

New Thread

HEA-NOC/20080410-5 [UPDATE] NUIG Network Transition Second Phase

HEA-NOC/20080410-5 [UPDATE] NUIG Network Transition Second Phase

Brian Nisbet

Thu, 26 Jun 2008 09:19:42 +0100

54 lines

HEA-NOC/20080410-5 [UPDATE] NUIG Network Transition Second Phase

Brian Nisbet

Mon, 23 Jun 2008 11:48:11 +0100

50 lines

New Thread

HEA-NOC/20080424-8 [CLOSE] Removing UCC Old Links

HEA-NOC/20080424-8 [CLOSE] Removing UCC Old Links

Cormac O'Cianain

Mon, 23 Jun 2008 16:12:33 +0100

37 lines

New Thread

HEA-NOC/20080513-1 [CLOSE] DCC fibre Cut

HEA-NOC/20080513-1 [CLOSE] DCC fibre Cut

Brian Nisbet

Tue, 10 Jun 2008 10:32:54 +0100

80 lines

New Thread

HEA-NOC/20080519-2 [CLOSE] NUI Migration to New IP Network

HEA-NOC/20080519-2 [CLOSE] NUI Migration to New IP Network

Garwin Liu

Tue, 3 Jun 2008 13:07:32 +0100

67 lines

New Thread

HEA-NOC/20080519-2 [UPDATE] NUI Migration to New IP Network

HEA-NOC/20080519-2 [UPDATE] NUI Migration to New IP Network

Garwin Liu

Tue, 3 Jun 2008 09:06:34 +0100

63 lines

New Thread

HEA-NOC/20080521-20 [CLOSE] Primary 10GE Link to Geant via London Down

HEA-NOC/20080521-20 [CLOSE] Primary 10GE Link to Geant via London Down

Garwin Liu

Tue, 3 Jun 2008 11:32:53 +0100

44 lines

New Thread

HEA-NOC/20080524-1 [CLOSE] GEANT Primary 10GE to London is down

HEA-NOC/20080524-1 [CLOSE] GEANT Primary 10GE to London is down

Gareth Eason

Tue, 3 Jun 2008 20:39:47 +0100

50 lines

New Thread

HEA-NOC/20080527-10 [CLOSE] Problems viewing GTI streams from RTE Streaming LAN

HEA-NOC/20080527-10 [CLOSE] Problems viewing GTI streams from RTE Streaming LAN

Orla McGann

Wed, 18 Jun 2008 18:33:53 +0100

57 lines

New Thread

HEA-NOC/20080527-12 [CLOSE] Bringing up second BGP session to CAO on cr2-kp

HEA-NOC/20080527-12 [CLOSE] Bringing up second BGP session to CAO on cr2-kp

Orla McGann

Thu, 5 Jun 2008 13:06:41 +0100

54 lines

New Thread

HEA-NOC/20080527-12 [UPDATE] Bringing up second BGP session to CAO on cr2-kp

HEA-NOC/20080527-12 [UPDATE] Bringing up second BGP session to CAO on cr2-kp

Orla McGann

Tue, 3 Jun 2008 15:59:29 +0100

48 lines

HEA-NOC/20080527-12 [UPDATE] Bringing up second BGP session to CAO on cr2-kp

Orla McGann

Tue, 3 Jun 2008 15:06:13 +0100

43 lines

New Thread

HEA-NOC/20080527-6 [CLOSE] WIT PoP Switch Outages

HEA-NOC/20080527-6 [CLOSE] WIT PoP Switch Outages

Cormac O'Cianain

Tue, 3 Jun 2008 10:21:21 +0100

37 lines

New Thread

HEA-NOC/20080527-9 [CLOSE] e|net Planned Work

HEA-NOC/20080527-9 [CLOSE] e|net Planned Work

Brian Nisbet

Wed, 4 Jun 2008 09:25:03 +0100

60 lines

New Thread

HEA-NOC/20080528-2 [UPDATE] Carlow College Migration to New IP Network

HEA-NOC/20080528-2 [UPDATE] Carlow College Migration to New IP Network

Garwin Liu

Wed, 4 Jun 2008 14:01:15 +0100

47 lines

New Thread

HEA-NOC/20080528-3 [UPDATE] Cessation of legacy Eircom ITnet links

HEA-NOC/20080528-3 [UPDATE] Cessation of legacy Eircom ITnet links

Owen Byrne

Fri, 20 Jun 2008 14:36:21 +0100

58 lines

HEA-NOC/20080528-3 [UPDATE] Cessation of legacy Eircom ITnet links

Owen Byrne

Thu, 19 Jun 2008 09:22:15 +0100

53 lines

HEA-NOC/20080528-3 [UPDATE] Cessation of legacy Eircom ITnet links

Owen Byrne

Wed, 4 Jun 2008 15:50:18 +0100

50 lines

New Thread

HEA-NOC/20080529-4 [CLOSE] Citywest Core Router Upgrade

HEA-NOC/20080529-4 [CLOSE] Citywest Core Router Upgrade

Brian Nisbet

Mon, 9 Jun 2008 11:21:52 +0100

48 lines

New Thread

HEA-NOC/20080529-4 [UPDATE] Citywest Core Router Upgrade

HEA-NOC/20080529-4 [UPDATE] Citywest Core Router Upgrade

Brian Nisbet

Thu, 5 Jun 2008 08:56:24 +0100

45 lines

New Thread

HEA-NOC/20080601-1 [CLOSE] Geant link to London

HEA-NOC/20080601-1 [CLOSE] Geant link to London

Gareth Eason

Tue, 10 Jun 2008 10:41:43 +0100

60 lines

New Thread

HEA-NOC/20080601-1 [UPDATE] Geant link to London

HEA-NOC/20080601-1 [UPDATE] Geant link to London

Gareth Eason

Tue, 3 Jun 2008 20:38:24 +0100

53 lines

HEA-NOC/20080601-1 [UPDATE] Geant link to London

Gareth Eason

Mon, 2 Jun 2008 16:44:48 +0100

46 lines

HEA-NOC/20080601-1 [UPDATE] Geant link to London

Rachael Holt

Sun, 1 Jun 2008 15:49:19 +0100

41 lines

New Thread

HEA-NOC/20080603-2 [OPEN] Planned maintenance in the NIRAN Pop in NWRC, Derry.

HEA-NOC/20080603-2 [OPEN] Planned maintenance in the NIRAN Pop in NWRC, Derry.

Orla McGann

Tue, 3 Jun 2008 14:29:25 +0100

41 lines

New Thread

HEA-NOC/20080603-3 [CLOSE] Setup bonded interface on samhain

HEA-NOC/20080603-3 [CLOSE] Setup bonded interface on samhain

Rob Gallagher

Tue, 10 Jun 2008 11:12:18 +0100

44 lines

New Thread

HEA-NOC/20080603-3 [OPEN] Setup bonded interface on samhain

HEA-NOC/20080603-3 [OPEN] Setup bonded interface on samhain

Rob Gallagher

Tue, 3 Jun 2008 16:45:11 +0100

36 lines

New Thread

HEA-NOC/20080603-3 [UPDATE] Setup bonded interface on samhain

HEA-NOC/20080603-3 [UPDATE] Setup bonded interface on samhain

Rob Gallagher

Wed, 4 Jun 2008 18:00:36 +0100

42 lines

HEA-NOC/20080603-3 [UPDATE] Setup bonded interface on samhain

Rob Gallagher

Wed, 4 Jun 2008 14:46:50 +0100

40 lines

New Thread

HEA-NOC/20080603-6 [CLOSE] Link Flaps on GEANT GENEVA Circuit

HEA-NOC/20080603-6 [CLOSE] Link Flaps on GEANT GENEVA Circuit

Paul Cosgrove

Thu, 19 Jun 2008 12:20:52 +0100

43 lines

New Thread

HEA-NOC/20080603-6 [OPEN] Link Flaps on GEANT GENEVA Circuit

HEA-NOC/20080603-6 [OPEN] Link Flaps on GEANT GENEVA Circuit

Paul Cosgrove

Tue, 3 Jun 2008 18:44:36 +0100

40 lines

New Thread

HEA-NOC/20080604-3 [CLOSE] Dublin and National Ring Fibre Moves

HEA-NOC/20080604-3 [CLOSE] Dublin and National Ring Fibre Moves

Brian Nisbet

Mon, 16 Jun 2008 09:51:49 +0100

50 lines

New Thread

HEA-NOC/20080604-3 [OPEN] Dublin and National Ring Fibre Moves

HEA-NOC/20080604-3 [OPEN] Dublin and National Ring Fibre Moves

Brian Nisbet

Wed, 4 Jun 2008 12:49:32 +0100

37 lines

New Thread

HEA-NOC/20080604-3 [UPDATE] Dublin and National Ring Fibre Moves

HEA-NOC/20080604-3 [UPDATE] Dublin and National Ring Fibre Moves

Garreth Malone

Thu, 12 Jun 2008 08:58:24 +0100

47 lines

HEA-NOC/20080604-3 [UPDATE] Dublin and National Ring Fibre Moves

Garreth Malone

Wed, 11 Jun 2008 09:14:51 +0100

44 lines

New Thread

HEA-NOC/20080604-7 [CLOSE] Carlow College Outage

HEA-NOC/20080604-7 [CLOSE] Carlow College Outage

Garwin Liu

Fri, 6 Jun 2008 11:20:17 +0100

32 lines

New Thread

HEA-NOC/20080604-7 [OPEN] Carlow College Outage

HEA-NOC/20080604-7 [OPEN] Carlow College Outage

Garwin Liu

Wed, 4 Jun 2008 19:02:17 +0100

29 lines

New Thread

HEA-NOC/20080605-6 [CLOSE] Mary Immaculate UL loss connectivity

HEA-NOC/20080605-6 [CLOSE] Mary Immaculate UL loss connectivity

Cormac O'Cianain

Thu, 5 Jun 2008 12:31:01 +0100

38 lines

New Thread

HEA-NOC/20080605-6 [OPEN] Mary Immaculate UL loss connectivity

HEA-NOC/20080605-6 [OPEN] Mary Immaculate UL loss connectivity

Cormac O'Cianain

Thu, 5 Jun 2008 11:55:10 +0100

28 lines

New Thread

HEA-NOC/20080605-6 [UPDATE] Mary Immaculate UL loss connectivity

HEA-NOC/20080605-6 [UPDATE] Mary Immaculate UL loss connectivity

Cormac O'Cianain

Thu, 5 Jun 2008 11:57:49 +0100

33 lines

New Thread

HEA-NOC/20080606-2 [CLOSE] Maintenace Work - Outage notice 18/06/2008

HEA-NOC/20080606-2 [CLOSE] Maintenace Work - Outage notice 18/06/2008

Garwin Liu

Wed, 18 Jun 2008 10:24:54 +0100

76 lines

New Thread

HEA-NOC/20080606-2 [OPEN] Maintenace Work - Outage notice 18/06/2008

HEA-NOC/20080606-2 [OPEN] Maintenace Work - Outage notice 18/06/2008

Garwin Liu

Fri, 6 Jun 2008 11:33:49 +0100

68 lines

New Thread

HEA-NOC/20080606-2 [UPDATE] Maintenace Work - Outage notice 18/06/2008

HEA-NOC/20080606-2 [UPDATE] Maintenace Work - Outage notice 18/06/2008

Garwin Liu

Fri, 6 Jun 2008 12:29:39 +0100

73 lines

New Thread

HEA-NOC/20080606-4 [CLOSE] Loss of connectivity Tallaght Hostpital

HEA-NOC/20080606-4 [CLOSE] Loss of connectivity Tallaght Hostpital

Cormac O'Cianain

Fri, 6 Jun 2008 14:57:41 +0100

32 lines

New Thread

HEA-NOC/20080606-4 [OPEN] Loss of connectivity Tallaght Hostpital

HEA-NOC/20080606-4 [OPEN] Loss of connectivity Tallaght Hostpital

Cormac O'Cianain

Fri, 6 Jun 2008 13:45:30 +0100

25 lines

New Thread

HEA-NOC/20080606-6 [OPEN] IT Tralee connection move

HEA-NOC/20080606-6 [OPEN] IT Tralee connection move

Dave Wilson

Fri, 6 Jun 2008 15:42:49 +0100

34 lines

New Thread

HEA-NOC/20080606-6 [UPDATE] IT Tralee connection move

HEA-NOC/20080606-6 [UPDATE] IT Tralee connection move

Dave Wilson

Thu, 12 Jun 2008 15:21:26 +0100

38 lines

New Thread

HEA-NOC/20080606-7 [OPEN] JANET connection recabling in Citywest

HEA-NOC/20080606-7 [OPEN] JANET connection recabling in Citywest

Dave Wilson

Fri, 6 Jun 2008 15:44:37 +0100

32 lines

New Thread

HEA-NOC/20080606-7 [UPDATE] JANET connection recabling in Citywest

HEA-NOC/20080606-7 [UPDATE] JANET connection recabling in Citywest

Dave Wilson

Thu, 12 Jun 2008 11:38:50 +0100

36 lines

New Thread

HEA-NOC/20080606-8 [CLOSE] DKIT Connectivity issues

HEA-NOC/20080606-8 [CLOSE] DKIT Connectivity issues

Kevin Dermody

Thu, 12 Jun 2008 13:09:45 +0100

45 lines

New Thread

HEA-NOC/20080606-8 [OPEN] DKIT Connectivity issues

HEA-NOC/20080606-8 [OPEN] DKIT Connectivity issues

Kevin Dermody

Fri, 6 Jun 2008 17:46:29 +0100

37 lines

New Thread

HEA-NOC/20080607-1 [CLOSE] Brief outage to St. Angela's College

HEA-NOC/20080607-1 [CLOSE] Brief outage to St. Angela's College

Tony Gray

Tue, 10 Jun 2008 16:29:32 +0100

51 lines

New Thread

HEA-NOC/20080607-1 [OPEN] Brief outage to St. Angela's College

HEA-NOC/20080607-1 [OPEN] Brief outage to St. Angela's College

Orla McGann

Sat, 7 Jun 2008 10:02:11 +0100

32 lines

New Thread

HEA-NOC/20080607-1 [UPDATE] Brief outage to St. Angela's College

HEA-NOC/20080607-1 [UPDATE] Brief outage to St. Angela's College

Tony Gray

Mon, 9 Jun 2008 16:01:57 +0100

48 lines

HEA-NOC/20080607-1 [UPDATE] Brief outage to St. Angela's College

Tony Gray

Mon, 9 Jun 2008 10:49:38 +0100

40 lines

HEA-NOC/20080607-1 [UPDATE] Brief outage to St. Angela's College

Orla McGann

Sat, 7 Jun 2008 16:34:59 +0100

39 lines

HEA-NOC/20080607-1 [UPDATE] Brief outage to St. Angela's College

Orla McGann

Sat, 7 Jun 2008 10:28:09 +0100

35 lines

New Thread

HEA-NOC/20080607-3 [CLOSE] Scheduled outage to Dark Fibre on NTFON

HEA-NOC/20080607-3 [CLOSE] Scheduled outage to Dark Fibre on NTFON

Aidan Carty

Mon, 16 Jun 2008 10:04:23 +0100

56 lines

New Thread

HEA-NOC/20080607-3 [OPEN] Scheduled outage to Dark Fibre on NTFON

HEA-NOC/20080607-3 [OPEN] Scheduled outage to Dark Fibre on NTFON

Orla McGann

Sat, 7 Jun 2008 11:07:40 +0100

53 lines

New Thread

HEA-NOC/20080607-4 [CLOSE] Alarms detected on BT circuit HEAN107523

HEA-NOC/20080607-4 [CLOSE] Alarms detected on BT circuit HEAN107523

Orla McGann

Sat, 7 Jun 2008 12:57:58 +0100

37 lines

New Thread

HEA-NOC/20080607-4 [OPEN] Alarms detected on BT circuit HEAN107523

HEA-NOC/20080607-4 [OPEN] Alarms detected on BT circuit HEAN107523

Orla McGann

Sat, 7 Jun 2008 12:25:48 +0100

27 lines

New Thread

HEA-NOC/20080609-5 [OPEN] IADT Down

HEA-NOC/20080609-5 [OPEN] IADT Down

Gareth Eason

Mon, 9 Jun 2008 15:56:14 +0100

30 lines

New Thread

HEA-NOC/20080609-5 [UPDATE] IADT Down

HEA-NOC/20080609-5 [UPDATE] IADT Down

Cillian Sharkey

Mon, 9 Jun 2008 18:13:13 +0100

39 lines

HEA-NOC/20080609-5 [UPDATE] IADT Down

Cillian Sharkey

Mon, 9 Jun 2008 18:01:59 +0100

37 lines

HEA-NOC/20080609-5 [UPDATE] IADT Down

Tony Gray

Mon, 9 Jun 2008 17:01:38 +0100

34 lines

New Thread

HEA-NOC/20080610-2 [OPEN] Routing policy change to General Internet

HEA-NOC/20080610-2 [OPEN] Routing policy change to General Internet

Daniel Lete

Tue, 10 Jun 2008 10:12:19 +0100

45 lines

New Thread

HEA-NOC/20080610-2 [UPDATE] Routing policy change to General Internet

HEA-NOC/20080610-2 [UPDATE] Routing policy change to General Internet

Daniel Lete

Wed, 18 Jun 2008 12:12:35 +0100

49 lines

New Thread

HEA-NOC/20080610-7 [CLOSE] restart of Primary DNS server ns.heanet.ie

HEA-NOC/20080610-7 [CLOSE] restart of Primary DNS server ns.heanet.ie

Aidan Carty

Wed, 11 Jun 2008 09:07:16 +0100

35 lines

New Thread

HEA-NOC/20080610-7 [OPEN] restart of Primary DNS server ns.heanet.ie

HEA-NOC/20080610-7 [OPEN] restart of Primary DNS server ns.heanet.ie

Aidan Carty

Tue, 10 Jun 2008 16:28:15 +0100

32 lines

New Thread

HEA-NOC/20080611-11 [CLOSE] Replacement of memory card in Drumline ADVA node.

HEA-NOC/20080611-11 [CLOSE] Replacement of memory card in Drumline ADVA node.

Brian Nisbet

Mon, 23 Jun 2008 13:07:23 +0100

44 lines

New Thread

HEA-NOC/20080611-11 [OPEN] Replacement of memory card in Drumline ADVA node.

HEA-NOC/20080611-11 [OPEN] Replacement of memory card in Drumline ADVA node.

Gareth Eason

Wed, 11 Jun 2008 12:01:03 +0100

37 lines

New Thread

HEA-NOC/20080611-15 [OPEN] An Chéim Migration

HEA-NOC/20080611-15 [OPEN] An Chéim Migration

Cormac O'Cianain

Wed, 11 Jun 2008 16:00:45 +0100

34 lines

New Thread

HEA-NOC/20080611-15 [UPDATE] An Chéim Migration

HEA-NOC/20080611-15 [UPDATE] An Chéim Migration

Cormac O'Cianain

Wed, 18 Jun 2008 09:34:03 +0100

42 lines

HEA-NOC/20080611-15 [UPDATE] An Chéim Migration

Cormac O'Cianain

Tue, 17 Jun 2008 17:35:06 +0100

39 lines

New Thread

HEA-NOC/20080611-16 [CLOSE] DC electrical work in CWT PoP

HEA-NOC/20080611-16 [CLOSE] DC electrical work in CWT PoP

Garreth Malone

Thu, 12 Jun 2008 18:11:30 +0100

48 lines

New Thread

HEA-NOC/20080611-16 [OPEN] DC electrical work in CWT PoP

HEA-NOC/20080611-16 [OPEN] DC electrical work in CWT PoP

Garreth Malone

Wed, 11 Jun 2008 16:36:57 +0100

42 lines

New Thread

HEA-NOC/20080611-16 [UPDATE] DC electrical work in CWT PoP

HEA-NOC/20080611-16 [UPDATE] DC electrical work in CWT PoP

Garreth Malone

Thu, 12 Jun 2008 09:21:23 +0100

45 lines

New Thread

HEA-NOC/20080612-15 [OPEN] e-Net Planned Outage Notice

HEA-NOC/20080612-15 [OPEN] e-Net Planned Outage Notice

Aidan Carty

Thu, 12 Jun 2008 17:21:19 +0100

32 lines

New Thread

HEA-NOC/20080612-15 [UPDATE] e-Net Planned Outage Notice

HEA-NOC/20080612-15 [UPDATE] e-Net Planned Outage Notice

Dave Wilson

Tue, 17 Jun 2008 13:00:06 +0100

36 lines

New Thread

HEA-NOC/20080612-4 [CLOSE] Removal of Old Redundant Connections to Old IP Network

HEA-NOC/20080612-4 [CLOSE] Removal of Old Redundant Connections to Old IP Network

Garwin Liu

Tue, 24 Jun 2008 11:46:42 +0100

36 lines

New Thread

HEA-NOC/20080612-4 [OPEN] Removal of Old Redundant Connections to Old IP Network

HEA-NOC/20080612-4 [OPEN] Removal of Old Redundant Connections to Old IP Network

Garwin Liu

Thu, 12 Jun 2008 09:38:18 +0100

33 lines

New Thread

HEA-NOC/20080613-1 [OPEN] Tipp Inst Site outage

HEA-NOC/20080613-1 [OPEN] Tipp Inst Site outage

Aidan Carty

Fri, 13 Jun 2008 09:06:22 +0100

24 lines

New Thread

HEA-NOC/20080613-1 [UPDATE] Tipp Inst Site outage

HEA-NOC/20080613-1 [UPDATE] Tipp Inst Site outage

Gareth Eason

Sat, 14 Jun 2008 10:42:42 +0100

62 lines

HEA-NOC/20080613-1 [UPDATE] Tipp Inst Site outage

Gareth Eason

Fri, 13 Jun 2008 16:46:27 +0100

56 lines

HEA-NOC/20080613-1 [UPDATE] Tipp Inst Site outage

Gareth Eason

Fri, 13 Jun 2008 15:56:29 +0100

47 lines

HEA-NOC/20080613-1 [UPDATE] Tipp Inst Site outage

Aidan Carty

Fri, 13 Jun 2008 14:11:55 +0100

37 lines

HEA-NOC/20080613-1 [UPDATE] Tipp Inst Site outage

Gareth Eason

Fri, 13 Jun 2008 13:21:34 +0100

34 lines

HEA-NOC/20080613-1 [UPDATE] Tipp Inst Site outage

Aidan Carty

Fri, 13 Jun 2008 10:32:44 +0100

28 lines

New Thread

HEA-NOC/20080613-11 [CLOSE] No connectivity to GTI

HEA-NOC/20080613-11 [CLOSE] No connectivity to GTI

Gareth Eason

Fri, 13 Jun 2008 15:37:34 +0100

59 lines

New Thread

HEA-NOC/20080613-11 [OPEN] No connectivity to GTI

HEA-NOC/20080613-11 [OPEN] No connectivity to GTI

Gareth Eason

Fri, 13 Jun 2008 12:52:03 +0100

35 lines

New Thread

HEA-NOC/20080613-11 [UPDATE] No connectivity to GTI

HEA-NOC/20080613-11 [UPDATE] No connectivity to GTI

Aidan Carty

Fri, 13 Jun 2008 14:30:17 +0100

47 lines

HEA-NOC/20080613-11 [UPDATE] No connectivity to GTI

Gareth Eason

Fri, 13 Jun 2008 13:38:24 +0100

45 lines

HEA-NOC/20080613-11 [UPDATE] No connectivity to GTI

Gareth Eason

Fri, 13 Jun 2008 12:57:39 +0100

38 lines

New Thread

HEA-NOC/20080613-17 [CLOSE] Repatching of Global Crossing fibre for GEANT in Esat-X

HEA-NOC/20080613-17 [CLOSE] Repatching of Global Crossing fibre for GEANT in Esat-X

Garreth Malone

Tue, 24 Jun 2008 15:31:16 +0100

45 lines

New Thread

HEA-NOC/20080613-17 [OPEN] Repatching of Global Crossing fibre for GEANT in Esat-X

HEA-NOC/20080613-17 [OPEN] Repatching of Global Crossing fibre for GEANT in Esat-X

Gareth Eason

Fri, 13 Jun 2008 18:14:08 +0100

30 lines

New Thread

HEA-NOC/20080613-17 [UPDATE] Repatching of Global Crossing fibre for GEANT in Esat-X

HEA-NOC/20080613-17 [UPDATE] Repatching of Global Crossing fibre for GEANT in Esat-X

Garreth Malone

Mon, 16 Jun 2008 22:35:14 +0100

43 lines

HEA-NOC/20080613-17 [UPDATE] Repatching of Global Crossing fibre for GEANT in Esat-X

Garreth Malone

Mon, 16 Jun 2008 21:12:04 +0100

37 lines

HEA-NOC/20080613-17 [UPDATE] Repatching of Global Crossing fibre for GEANT in Esat-X

Garreth Malone

Mon, 16 Jun 2008 12:18:17 +0100

34 lines

New Thread

HEA-NOC/20080613-9 [OPEN/CLOSE] Link to GTI flapped

HEA-NOC/20080613-9 [OPEN/CLOSE] Link to GTI flapped

Gareth Eason

Fri, 13 Jun 2008 12:04:04 +0100

32 lines

New Thread

HEA-NOC/20080616-2 [OPEN] ESB fibre work affecting Galway PoP

HEA-NOC/20080616-2 [OPEN] ESB fibre work affecting Galway PoP

Dave Wilson

Mon, 16 Jun 2008 12:31:13 +0100

33 lines

New Thread

HEA-NOC/20080616-4 [CLOSE] HEAnet NOC Phone System At Risk

HEA-NOC/20080616-4 [CLOSE] HEAnet NOC Phone System At Risk

Brian Nisbet

Tue, 17 Jun 2008 13:18:46 +0100

39 lines

New Thread

HEA-NOC/20080616-4 [OPEN] HEAnet NOC Phone System At Risk

HEA-NOC/20080616-4 [OPEN] HEAnet NOC Phone System At Risk

Brian Nisbet

Mon, 16 Jun 2008 16:04:58 +0100

36 lines

New Thread

HEA-NOC/20080617-6 [OPEN] Multicast routing not working at IT Athlone

HEA-NOC/20080617-6 [OPEN] Multicast routing not working at IT Athlone

Paul Cosgrove

Tue, 17 Jun 2008 12:33:20 +0100

33 lines

New Thread

HEA-NOC/20080617-8 [OPEN] TCD transition to new IP network

HEA-NOC/20080617-8 [OPEN] TCD transition to new IP network

Dave Wilson

Tue, 17 Jun 2008 17:27:26 +0100

36 lines

New Thread

HEA-NOC/20080617-8 [UPDATE] TCD transition to new IP network

HEA-NOC/20080617-8 [UPDATE] TCD transition to new IP network

Dave Wilson

Fri, 27 Jun 2008 16:10:18 +0100

46 lines

HEA-NOC/20080617-8 [UPDATE] TCD transition to new IP network

Dave Wilson

Thu, 26 Jun 2008 09:22:06 +0100

43 lines

HEA-NOC/20080617-8 [UPDATE] TCD transition to new IP network

Dave Wilson

Thu, 26 Jun 2008 08:07:49 +0100

39 lines

New Thread

HEA-NOC/20080618-6 [CLOSE] Migration of ns.heanet.ie to new hardware

HEA-NOC/20080618-6 [CLOSE] Migration of ns.heanet.ie to new hardware

Rob Gallagher

Mon, 30 Jun 2008 10:53:17 +0100

48 lines

New Thread

HEA-NOC/20080618-6 [OPEN] Migration of ns.heanet.ie to new hardware

HEA-NOC/20080618-6 [OPEN] Migration of ns.heanet.ie to new hardware

Rob Gallagher

Wed, 18 Jun 2008 12:22:27 +0100

37 lines

New Thread

HEA-NOC/20080618-6 [UPDATE] Migration of ns.heanet.ie to new hardware

HEA-NOC/20080618-6 [UPDATE] Migration of ns.heanet.ie to new hardware

Rob Gallagher

Wed, 25 Jun 2008 08:26:53 +0100

46 lines

HEA-NOC/20080618-6 [UPDATE] Migration of ns.heanet.ie to new hardware

Rob Gallagher

Tue, 24 Jun 2008 16:30:49 +0100

43 lines

New Thread

HEA-NOC/20080619-5 [CLOSE] IPv6 connectivity to WIT disrupted

HEA-NOC/20080619-5 [CLOSE] IPv6 connectivity to WIT disrupted

Rob Gallagher

Fri, 20 Jun 2008 09:42:14 +0100

39 lines

New Thread

HEA-NOC/20080619-5 [OPEN] IPv6 connectivity to WIT disrupted

HEA-NOC/20080619-5 [OPEN] IPv6 connectivity to WIT disrupted

Rob Gallagher

Thu, 19 Jun 2008 14:45:16 +0100

32 lines

New Thread

HEA-NOC/20080619-5 [UPDATE] IPv6 connectivity to WIT disrupted

HEA-NOC/20080619-5 [UPDATE] IPv6 connectivity to WIT disrupted

Rob Gallagher

Thu, 19 Jun 2008 15:57:59 +0100

35 lines

New Thread

HEA-NOC/20080621-1 [OPEN] The circuit to the Hermes Project schools is down

HEA-NOC/20080621-1 [OPEN] The circuit to the Hermes Project schools is down

Brian Boyle

Sat, 21 Jun 2008 15:06:08 +0100

26 lines

New Thread

HEA-NOC/20080621-1 [UPDATE] The circuit to the Hermes Project schools is down

HEA-NOC/20080621-1 [UPDATE] The circuit to the Hermes Project schools is down

Brian Boyle

Sat, 21 Jun 2008 19:20:25 +0100

30 lines

New Thread

HEA-NOC/20080623-11 [CLOSE] UCD IP Link Transition

HEA-NOC/20080623-11 [CLOSE] UCD IP Link Transition

Brian Nisbet

Thu, 26 Jun 2008 09:18:48 +0100

47 lines

New Thread

HEA-NOC/20080623-11 [OPEN] UCD IP Link Transition

HEA-NOC/20080623-11 [OPEN] UCD IP Link Transition

Brian Nisbet

Mon, 23 Jun 2008 15:06:12 +0100

40 lines

New Thread

HEA-NOC/20080623-11 [UPDATE] UCD IP Link Transition

HEA-NOC/20080623-11 [UPDATE] UCD IP Link Transition

Brian Nisbet

Tue, 24 Jun 2008 18:47:41 +0100

45 lines

New Thread

HEA-NOC/20080623-13 [OPEN] Removal of Old Circuits

HEA-NOC/20080623-13 [OPEN] Removal of Old Circuits

Cormac O'Cianain

Mon, 23 Jun 2008 18:13:29 +0100

28 lines

New Thread

HEA-NOC/20080623-13 [UPDATE] Removal of Old Circuits

HEA-NOC/20080623-13 [UPDATE] Removal of Old Circuits

Cormac O'Cianain

Wed, 25 Jun 2008 18:32:34 +0100

31 lines

New Thread

HEA-NOC/20080623-9 [CLOSE] Repair work on fibres in the HEAnet Servecentric PoP

HEA-NOC/20080623-9 [CLOSE] Repair work on fibres in the HEAnet Servecentric PoP

Orla McGann

Tue, 24 Jun 2008 11:28:44 +0100

64 lines

New Thread

HEA-NOC/20080623-9 [OPEN] Repair work on fibres in the HEAnet Servecentric PoP

HEA-NOC/20080623-9 [OPEN] Repair work on fibres in the HEAnet Servecentric PoP

Paul Cosgrove

Mon, 23 Jun 2008 13:21:24 +0100

40 lines

New Thread

HEA-NOC/20080623-9 [UPDATE] Repair work on fibres in the HEAnet Servecentric PoP

HEA-NOC/20080623-9 [UPDATE] Repair work on fibres in the HEAnet Servecentric PoP

Orla McGann

Tue, 24 Jun 2008 07:56:31 +0100

60 lines

HEA-NOC/20080623-9 [UPDATE] Repair work on fibres in the HEAnet Servecentric PoP

Paul Cosgrove

Mon, 23 Jun 2008 16:53:59 +0100

53 lines

HEA-NOC/20080623-9 [UPDATE] Repair work on fibres in the HEAnet Servecentric PoP

Paul Cosgrove

Mon, 23 Jun 2008 16:01:38 +0100

46 lines

New Thread

HEA-NOC/20080624-1 [OPEN] An Cheim Failover Testing

HEA-NOC/20080624-1 [OPEN] An Cheim Failover Testing

Owen Byrne

Tue, 24 Jun 2008 08:49:58 +0100

29 lines

New Thread

HEA-NOC/20080624-1 [UPDATE] An Cheim Failover Testing

HEA-NOC/20080624-1 [UPDATE] An Cheim Failover Testing

Garwin Liu

Tue, 24 Jun 2008 13:08:22 +0100

34 lines

New Thread

HEA-NOC/20080624-12 [OPEN] Luas Works affecting NCI backup circuit

HEA-NOC/20080624-12 [OPEN] Luas Works affecting NCI backup circuit

Paul Cosgrove

Tue, 24 Jun 2008 15:50:05 +0100

44 lines

New Thread

HEA-NOC/20080624-13 [OPEN] ISDN gateway service line maintenance

HEA-NOC/20080624-13 [OPEN] ISDN gateway service line maintenance

Justin Hourigan

Tue, 24 Jun 2008 15:52:26 +0100

31 lines

New Thread

HEA-NOC/20080624-3 [CLOSE] Network maintenance at IT Tralee

HEA-NOC/20080624-3 [CLOSE] Network maintenance at IT Tralee

Garwin Liu

Mon, 30 Jun 2008 11:45:16 +0100

34 lines

New Thread

HEA-NOC/20080624-3 [OPEN] Network maintenance at IT Tralee

HEA-NOC/20080624-3 [OPEN] Network maintenance at IT Tralee

Paul Cosgrove

Tue, 24 Jun 2008 09:53:59 +0100

28 lines

New Thread

HEA-NOC/20080624-5 [CLOSE] Netflow Server Reboot Required

HEA-NOC/20080624-5 [CLOSE] Netflow Server Reboot Required

Cormac O'Cianain

Tue, 24 Jun 2008 17:47:13 +0100

36 lines

New Thread

HEA-NOC/20080624-5 [OPEN] Netflow Server Reboot Required

HEA-NOC/20080624-5 [OPEN] Netflow Server Reboot Required

Cormac O'Cianain

Tue, 24 Jun 2008 11:15:37 +0100

29 lines

New Thread

HEA-NOC/20080624-7 [CLOSE] Serial backup circuit to NCRI down

HEA-NOC/20080624-7 [CLOSE] Serial backup circuit to NCRI down

Paul Cosgrove

Tue, 24 Jun 2008 11:57:31 +0100

33 lines

New Thread

HEA-NOC/20080624-7 [OPEN] Serial backup circuit to NCRI down

HEA-NOC/20080624-7 [OPEN] Serial backup circuit to NCRI down

Paul Cosgrove

Tue, 24 Jun 2008 11:45:39 +0100

29 lines

New Thread

HEA-NOC/20080624-9 [OPEN] Scheduled Maintenance on JANET network 8th July 2008

HEA-NOC/20080624-9 [OPEN] Scheduled Maintenance on JANET network 8th July 2008

Paul Cosgrove

Tue, 24 Jun 2008 12:51:27 +0100

40 lines

New Thread

HEA-NOC/20080624-9 [UPDATE] Scheduled Maintenance on JANET network 8th July 2008

HEA-NOC/20080624-9 [UPDATE] Scheduled Maintenance on JANET network 8th July 2008

Paul Cosgrove

Tue, 24 Jun 2008 12:57:40 +0100

49 lines

New Thread

HEA-NOC/20080626-8 [CLOSE] UL wish to cease using the hotstandby service.

HEA-NOC/20080626-8 [CLOSE] UL wish to cease using the hotstandby service.

Paul Cosgrove

Fri, 27 Jun 2008 09:54:58 +0100

55 lines

New Thread

HEA-NOC/20080626-8 [OPEN] UL wish to cease using the hotstandby service.

HEA-NOC/20080626-8 [OPEN] UL wish to cease using the hotstandby service.

Paul Cosgrove

Thu, 26 Jun 2008 13:41:32 +0100

41 lines

New Thread

HEA-NOC/20080626-8 [UPDATE] UL wish to cease using the hotstandby service.

HEA-NOC/20080626-8 [UPDATE] UL wish to cease using the hotstandby service.

Paul Cosgrove

Thu, 26 Jun 2008 15:31:34 +0100

52 lines

New Thread

HEA-NOC/20080627-1 [OPEN] IRCSET migration within HEA IT infrastructure.

HEA-NOC/20080627-1 [OPEN] IRCSET migration within HEA IT infrastructure.

Paul Cosgrove

Fri, 27 Jun 2008 10:23:03 +0100

39 lines

New Thread

HEA-NOC/20080627-1 [UPDATE] IRCSET migration within HEA IT infrastructure.

HEA-NOC/20080627-1 [UPDATE] IRCSET migration within HEA IT infrastructure.

Paul Cosgrove

Fri, 27 Jun 2008 16:17:06 +0100

56 lines

HEA-NOC/20080627-1 [UPDATE] IRCSET migration within HEA IT infrastructure.

Paul Cosgrove

Fri, 27 Jun 2008 15:44:31 +0100

52 lines

New Thread

HEA-NOC/20080630-1 [OPEN] BT Fibre Maintenance - Cork <-> Mallow 20080710 - 23:50

HEA-NOC/20080630-1 [OPEN] BT Fibre Maintenance - Cork <-> Mallow 20080710 - 23:50

Garwin Liu

Mon, 30 Jun 2008 10:06:22 +0100

35 lines

Advanced Options


Options

Log In

Log In

Get Password

Get Password


Search Archives

Search Archives


Subscribe or Unsubscribe

Subscribe or Unsubscribe


Archives

November 2017
October 2017
September 2017
August 2017
July 2017
June 2017
May 2017
April 2017
March 2017
February 2017
January 2017
December 2016
November 2016
October 2016
September 2016
August 2016
July 2016
June 2016
May 2016
April 2016
February 2016
January 2016
December 2015
November 2015
October 2015
September 2015
August 2015
July 2015
June 2015
May 2015
April 2015
March 2015
February 2015
January 2015
December 2014
November 2014
October 2014
September 2014
August 2014
July 2014
June 2014
May 2014
April 2014
March 2014
February 2014
January 2014
December 2013
November 2013
October 2013
September 2013
August 2013
July 2013
June 2013
May 2013
April 2013
March 2013
February 2013
January 2013
December 2012
November 2012
October 2012
September 2012
August 2012
July 2012
June 2012
May 2012
April 2012
March 2012
February 2012
January 2012
December 2011
November 2011
October 2011
September 2011
August 2011
July 2011
June 2011
May 2011
April 2011
March 2011
February 2011
January 2011
December 2010
November 2010
October 2010
September 2010
August 2010
July 2010
June 2010
May 2010
April 2010
March 2010
February 2010
January 2010
December 2009
November 2009
October 2009
September 2009
August 2009
July 2009
June 2009
May 2009
April 2009
March 2009
February 2009
January 2009
December 2008
November 2008
October 2008
September 2008
August 2008
July 2008
June 2008
May 2008
April 2008
March 2008
February 2008
January 2008
December 2007
November 2007
October 2007
September 2007
August 2007
July 2007
June 2007
May 2007
April 2007
March 2007
February 2007
January 2007
December 2006
November 2006
October 2006
September 2006
August 2006
July 2006
June 2006
May 2006
April 2006
March 2006
February 2006
January 2006
December 2005
November 2005
October 2005
September 2005
August 2005
July 2005
June 2005
May 2005
April 2005
March 2005
February 2005
January 2005
December 2004
November 2004
October 2004
September 2004
August 2004
July 2004
June 2004
May 2004
April 2004
March 2004
February 2004
January 2004
December 2003
November 2003
October 2003
September 2003
August 2003
July 2003
June 2003
May 2003
April 2003
March 2003
February 2003
January 2003
December 2002
November 2002
October 2002
September 2002
August 2002
July 2002
June 2002
May 2002
April 2002
March 2002
February 2002
January 2002
December 2001
November 2001
October 2001
September 2001
August 2001
July 2001
June 2001
May 2001
April 2001
March 2001
February 2001
January 2001
December 2000
November 2000
October 2000
September 2000
August 2000
July 2000
June 2000
May 2000
April 2000
March 2000
February 2000
January 2000
December 1999
November 1999
October 1999
September 1999
August 1999
July 1999
June 1999
May 1999
April 1999
March 1999
February 1999
January 1999
December 1998
November 1998
October 1998
September 1998
August 1998
July 1998
June 1998
May 1998
April 1998
March 1998
February 1998
January 1998
December 1997
November 1997
October 1997
September 1997
August 1997
July 1997
June 1997
May 1997
April 1997
March 1997
February 1997
January 1997
December 1996
November 1996
October 1996
September 1996
August 1996
July 1996
June 1996
May 1996
April 1996
March 1996
February 1996
January 1996
December 1995
November 1995
October 1995
September 1995
August 1995
July 1995
June 1995
May 1995
April 1995
March 1995
February 1995
January 1995
December 1994
November 1994
October 1994
September 1994
August 1994
July 1994
June 1994
May 1994
April 1994
March 1994
February 1994
January 1994
December 1993
November 1993
October 1993
September 1993
August 1993
July 1993
June 1993

ATOM RSS1 RSS2



LISTSERV.HEANET.IE

Secured by F-Secure Anti-Virus CataList Email List Search Powered by the LISTSERV Email List Manager